The second angle of a building lot is three times as large as the first,The third angle is 10 degrees more than the sum of the other two angles.What is the third angle?

Answer by checkley79(3341)

SECOND ANGLE=3F
THIRD ANGLE=F+S+10
F+S+T=180
F+3F+(F+S+10)=180
4F+F+(3F)+10=180
8F=180-10
8F=170
F=170/8
F=21.25 DEGREES FOR THE FIRST ANGLE.
3*21.25=63.75 DEGREES FOR THE SECOND ANGLE.
21.25+63.75+10=95 DEGREES FOR THE THIRD ANGLE.
21.25+63.75+95=180
180=180